OpenStudy (parissfrance09):

If the diameter of the circle above is 4 mm, what is the circumference of the circle? A.16 mm B.2 mm C.4 mm D.8 mm

Nnesha (nnesha):

\[\huge\rm C = \pi \times d\] if you have diameter use thsi formula to find C

OpenStudy (parissfrance09):

i got 12.56???

Nnesha (nnesha):

and none of them is right answer :D

OpenStudy (parissfrance09):

im confused these were the only choices

Nnesha (nnesha):

wait

OpenStudy (texaschic101):

I got 12.56 as well

OpenStudy (parissfrance09):

so which one should i choose

OpenStudy (wormynut):



OpenStudy (texaschic101):

you shouldn't....they are all wrong...tell your teacher. Show the work you did and the answer you got, which was 12.56...thats what I would do

OpenStudy (alexandervonhumboldt2):

they forgot to write pi i think

OpenStudy (parissfrance09):

the problem is im homeschooled and these are online classes

Nnesha (nnesha):

yep 2 pi r is a formula but when you divide diameter and then multiply by 2pi you should get same answer so 12.57

OpenStudy (wormynut):



OpenStudy (parissfrance09):

so the answer is 2mm??

Nnesha (nnesha):

it can't be 2

OpenStudy (texaschic101):

could the answer be C and they just forgot to put pi ?

OpenStudy (parissfrance09):

maybe just tell me one to choose

OpenStudy (alexandervonhumboldt2):

circumfence=diameter*pi or 2*radius*pi

Nnesha (nnesha):

\[\huge\rm C = \pi \times d \] diameter is 4mm so replace d by 4 i don't if you have to use pi or not bec none of the answer choice have pi

OpenStudy (alexandervonhumboldt2):

i tink C if guessing they forgot to write pi

OpenStudy (texaschic101):

lol...I would put C and then write multiplied by pi (3.14)

OpenStudy (alexandervonhumboldt2):

maybe it didn't copied

OpenStudy (alexandervonhumboldt2):

C

OpenStudy (parissfrance09):

okay hold on ima pick c

OpenStudy (alexandervonhumboldt2):

that C for sure @parissfrance09

Nnesha (nnesha):

you can't say for sure

OpenStudy (alexandervonhumboldt2):

i know for sure they forgot pi

OpenStudy (parissfrance09):

c was right it was supposed to say 4 times pie the pie didnt show up

OpenStudy (alexandervonhumboldt2):

rthus for sure it is C

OpenStudy (alexandervonhumboldt2):

ok i'm so smartr

OpenStudy (texaschic101):

ok...now I know I am not going crazy

OpenStudy (alexandervonhumboldt2):

haaha

Nnesha (nnesha):

that's not a little mistake \[C = \pi \times d\] is not equal to 4mm